Using laws of exponents, simplify and write the answer in exponential form: 7^x  × 7^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of multiplying the powers with same base = a^m× a^n= a^m + n
Then,
= (7)^x +2

Thanks